Find the foci of the hyperbola with the equation (y2 / 4) – (x2 / 49) = 1
Answer by nyc_function(2741)

Your hyperbola is of the form y^2/b^2 - x^2/a^2 = 1.
We need to find a and b to start with.
b^2 = 4
sqrt{b^2} = sqrt{4}
b = 2
a^2 = 49
sqrt{a^2} = sqrt{49}
a = 7
We now know that a = 7 and b = 2.
The foci will be points (-c,0) and (c,0). This means we need to find c.
We can c using c = sqrt{a^2 + b^2}.
c = sqrt{7^2 + 2^2}
c = sqrt{49 + 4}
c = sqrt{53}
The foci are (-sqrt{53}, 0) and (sqrt{53}, 0).
Did you follow?